OpenStudy (anonymous):

Angle 1 and angle X are complementary. Angle 2 and angle X are complementary. Angle 1=2n+2 and angle 2= n + 32 Find n.

OpenStudy (anonymous):

2n+2+n+32=90 Add like terms: 3n+34=90 Subtract 34 from 90 3n=56 Divide 3 by 56.. 56 --- 3 You get a fraction. = 56/3